OpenStudy (anonymous):

Use a calculator to find the measure of acute angle theta to the nearest tenth of a degree.

OpenStudy (anonymous):

\[\cot \theta=7.596\]

OpenStudy (anonymous):

@douglaswinslowcooper @iPwnBunnies @Mason_Jones573 @mathmale @rock_mit182 @twvogels

OpenStudy (anonymous):

@geerky42

OpenStudy (ipwnbunnies):

Cotangent is the same as 1/tangent.

OpenStudy (ipwnbunnies):

\[\frac{1}{\tan \theta} = 7.596\]

OpenStudy (ipwnbunnies):

Solve for tan(theta). Then, use the arctangent function on your calculator.

OpenStudy (anonymous):

thanks @iPwnBunnies ! I figured it out I have another question, I will tag you.

OpenStudy (ipwnbunnies):

Ok.